OpenStudy (anonymous):

How many positive and negative real zeros does P(x) = x^5 + 4x^3 - x^2 + 6x have?

OpenStudy (softballgirl372015):

you have to use descartes rules of signs. did you learn about that yet?

Directrix (directrix):

From x^5 + 4x^3 - x^2 + 6x, factor out the common monomial term of x. x^5 + 4x^3 - x^2 + 6x = x*(? + ? +?+ ?)

Directrix (directrix):

@learnfast x^5 + 4x^3 - x^2 + 6x = x* (x^4 + ? +?+ ?)